Direct Method and Overhead Rates

Jasmine Company manufactures both pesticide and liquid fertilizer, with each product manufactured in separate departments. Three support departments support the production departments: Power, General Factory, and Purchasing. Budgeted data on the five departments are as follows:


Support Departments
Producing Departments
Power General
Factory
Purchasing Pesticide Liquid
Fertilizer
Overhead $80,000 $316,000    $165,000    $78,500 $107,400
Square feet 1,500 —    1,500    4,200 4,800
Machine hours 1,403    1,345    24,000 8,000
Purchase orders 20    40    7    120 60

The company does not break overhead into fixed and variable components. The bases for allocation are power—machine hours; general factory—square feet; and purchasing—purchase orders.

Required:

1. Allocate the overhead costs to the producing departments using the direct method. If required, round your allocation ratios to four decimal places and round allocated costs to the nearest dollar and use the rounded values for the subsequent calculations.

Cost assignment:

Pesticide Liquid Fertilizer
Direct costs $ $
Power
General Factory
Purchasing
Total $ $

2. Using machine hours, compute departmental overhead rates. (Round the overhead rates to the nearest cent.)

Departmental overhead rates
Pesticide $ per machine hour
Liquid Fertilizer $ per machine hour
Answer #1

Answers

  • Requirement 1

Pesticide

Liquid Fertilizer

Direct costs

$78,500

$107,400

Power

$60,000

$20,000

General Factory

$147,477

$168,523

Purchasing

$110,006

$54,995

Total

$395,983

$350,917

--Working

Pesticide

Liquid Fertilizer

Direct costs

=+F15968

=+G15968

Power

=80000*0.75

=80000*0.25

General Factory

=316000*0.4667

=316000*0.5333

Purchasing

=165000*0.6667

=165000*0.3333

>ratios

Pesticide

Liquid

Total

Square feet

4,200

4,800

9,000

Ratio for allocation

0.4667

0.5333

Machine hours

24,000

8,000

32,000

Ratio for allocation

0.7500

0.2500

Purchase orders

120

60

180

Ratio for allocation

0.6667

0.3333

  • Requirement 2

Departmental Overhead rate

Pesticide

$                                16.50

per machine hour

Liquid Fertilizer

$                                43.86

per machine hour

--Working

Total Cost

Total machine hours

Departmental Overhead rate

Pesticide

$395,983

24,000

$                                16.50

per machine hour

Liquid Fertilizer

$350,917

8,000

$                                43.86

per machine hour